# Bound excited states of Fröhlich polarons in one dimension

### v1

@@ Abstract
As seen in Fig. 2, the first bound state with excitation energy below $1\hbar\omega_{\mathrm{LO}}$ appears at $\alpha\approx 1.73$, in the intermediate coupling regime. While the excitation energies in the discretized phonon continuum strongly depend on the box size $L$, this dependence disappears for the excited state below the continuum threshold. The value of $\alpha$ where the bound state emerges was found by repeating calculations with different $L$ at $\alpha=1.7$ and $\alpha=1.8$, which resulted in the consistent crossing point shown in the inset of Fig. 2. For $L=10l_{0}$, the threshold was found to be $\alpha=1.731\pm 0.003$.
The excitation energies were found to converge more slowly in $L$ than the ground state energy, showing that the excited states are larger in real space than the ground state. However, the excitation energies converge more quickly in the momentum cutoff $k_{c}$ than the absolute energies, and $k_{c}\approx 4\pi/l_{0}$ is large enough to find accurate excitation energies in the range of $\alpha$ reported here. See the Supplemental Material for more details [50].
Properties of bound excited states – The properties of the eigenstates can be studied more closely by looking at the density of virtual phonons in each mode, as shown in Fig. 3. The ground state gains more phonons as the coupling strength increases, with a larger density in the lower momentum modes. The continuum threshold state $|\mathrm{C}\rangle$ has increased phonon density only in the $k=0$ mode, as expected.
Figure 3: The density of virtual phonons in each mode $\langle j|a_{k}^{\dagger}a_{k}|j\rangle$ for the ground state $|j\rangle=|\mathrm{GS}\rangle$ and first two excited states, with $L=6l_{0}$ and $k_{c}=4\pi/l_{0}$, for various coupling strengths. The state $|\mathrm{C}\rangle$ is at the start of the continuum, and always has an excitation energy of $1\hbar\omega_{\mathrm{LO}}$. The state $|\mathrm{B}\rangle$ is a bound excited state for $\alpha\gtrsim 1.73$.
[…]
Spectral weight – The spectral weight of the state $|j\rangle$ at the total momentum $P$ from the Hamiltonian (Bound excited states of Fröhlich polarons in one dimension)
@@
Figure 4: Spectral weight $Z_{j}^{(0)}$ of the three lowest energy states vs. the coupling strength $\alpha$ for $k_{c}=4\pi/l_{0}$ and two different box sizes. For $\alpha>1.73$ marked by the vertical dashed line the state $|\mathrm{B}\rangle$ becomes a bound excited state. The dotted line is the first order perturbation theory result $21-\alpha/$ 2 for the ground state.
Outlook – An outstanding question concerns the relevance of our findings to materials [5] or experimentally accessible synthetic quantum systems [61, 62]. The closest related candidate systems may be narrow quantum wires with strong surface optical phonons [63]. Recently predicted discrete excited phonon branches near a Mott-insulator–superfluid phase transition in a Bose-Hubbard model [64] may be a related phenomenon to the excited states studied in this work. An interesting avenue of future research will be to extend the FCIQMC approach to the three-dimensional Fröhlich model to find the threshold coupling constant for the bound excited state and independently check the diagrammatic Monte Carlo results of Ref. [8], which indicated the absence of bound excited states up to values of $\alpha\approx 8$. It will further be valuable to understand the class of problems in which the resource requirements for sign-problem amelioration by walker annihilation in projector quantum Monte Carlo decouple from the dimension of Hilbert space.
@@ Acknowledgements.
Acknowledgements – We are grateful to Dmytro Kolisnyk for his help in working out the spectrum of the Hessian in the End Matter. This work was supported by the Marsden Fund of New Zealand (contract no. MAU2007) from government funding administered by the Royal Society Te Apārangi and by a summer scholarship from Te Whai Ao – Dodd-Walls Centre for Photonic and Quantum Technologies and the Physics Department, University of Auckland. We acknowledge support by the New Zealand eScience Infrastructure (NeSI) high-performance computing facilities in the form of a merit project allocation. Data availability – All numerical calculations were performed using the open source software Rimu.jl [27]. The data that support the findings of this article are openly available [65].

### v2

@@ VII Excited state energies
As seen in Fig. 2, the first bound state with excitation energy below $1\hbar\omega_{\mathrm{LO}}$ appears at $\alpha\approx 1.73$, in the intermediate coupling regime. While the excitation energies in the discretized phonon continuum strongly depend on the box size $L$, this dependence disappears for the excited state below the continuum threshold. The value of $\alpha$ where the first, even-parity bound state emerges was found by repeating calculations with different $L$ at $\alpha=1.7$ and $\alpha=1.8$, which resulted in the consistent interpolated crossing point shown in the inset of Fig. 2. For $L=10l_{0}$, the threshold was found to be $\alpha=1.731\pm 0.003$. The second bound state is of odd parity in the strong coupling limit, but not seen to breach the continuum threshold in the available numerical data. It is expected to appear at stronger coupling.
The excitation energies were found to converge more slowly in $L$ than the ground state energy, showing that the excited states are larger in real space than the ground state. However, the excitation energies converge more quickly in the momentum cutoff $k_{c}$ than the absolute energies, and $k_{c}=4\pi/l_{0}$ is large enough to find accurate excitation energies in the range of $\alpha$ reported here. For the bound excited state, the energy is converged at $L=6l_{0}$ and $k_{c}=4\pi/l_{0}$. See Appendix B for more details.
VIII Spectral weight
The spectral weight of the state $|j\rangle$ at the total momentum $P$ from the Hamiltonian (IV)
@@ VIII Spectral weight
Figure 3: Spectral weight $Z_{j}^{(0)}$ of the three lowest energy states vs. the coupling strength $\alpha$ for $k_{c}=4\pi/l_{0}$ and two different box sizes. For $\alpha>1.73$ marked by the vertical dashed line the state $|\mathrm{B}\rangle$ becomes a bound excited state. The dotted line is the first order perturbation theory result $1-\alpha/2$ for the ground state.
The number of phonons of the discrete states, $N_{|\mathrm{GS}\rangle}\approx 1.94$ and $N_{|\mathrm{B}\rangle}\approx 3.68$ at $\alpha=2$ (and $L=6l_{0}$, $k_{c}=4\pi/l_{0}$), reveals that extra phonons are bound by the polaron in the first excited state. This number is larger than in the strong coupling limit $\alpha\to\infty$ where $N_{|\mathrm{B}\rangle}-N_{|\mathrm{GS}\rangle}=0.6475$. Plots of the phonon densities are shown in Appendix D.
IX Outlook
An outstanding question concerns the relevance of our findings to materials [5] or experimentally accessible synthetic quantum systems [67, 68]. Polaron effects are expected to be relevant in various quasi-one-dimensional settings, such as carbon nanotubes [34, 69, 39, 40], quantum wires [70], or polymer chains [5]. Being a foundational and highly simplified model, however, the Fröhlich polaron model typically has to be extended to be applied to realistic materials [5]. Recently predicted discrete excited phonon branches near a Mott-insulator–superfluid phase transition in a Bose-Hubbard model [71] may be a related phenomenon to the excited states studied in this work. An interesting avenue of future research will be to extend the FCIQMC approach to the three-dimensional Fröhlich model to find the threshold coupling constant for the bound excited state and independently check the diagrammatic Monte Carlo results of Ref. [8], which indicated the absence of bound excited states up to values of $\alpha\approx 8$. The question remains as to the threshold for the appearance of the second, and further, bound excited states. Due to the large computational cost of large-$\alpha$ simulations, approaches such as importance sampling may be needed to improve the efficiency of the method in order to find these thresholds. It will further be valuable to understand the class of problems in which the resource requirements for sign-problem amelioration by walker annihilation in projector quantum Monte Carlo decouple from the dimension of Hilbert space.
@@ Acknowledgements.
We are grateful to Dmytro Kolisnyk for his help in working out the spectrum of the Hessian. This work was supported by the Marsden Fund of New Zealand (contract no. MAU2007) from government funding administered by the Royal Society Te Apārangi and by a summer scholarship from Te Whai Ao – Dodd-Walls Centre for Photonic and Quantum Technologies and the Physics Department, University of Auckland. We acknowledge support by the New Zealand eScience Infrastructure (NeSI) high-performance computing facilities in the form of a merit project allocation.
Data availability
All numerical calculations were performed using the open source software Rimu.jl [29]. The data that support the findings of this article are openly available [72].
Appendix A Ground State Energy Convergence
The energy of the ground state calculated using our computational model has a systematic discretization error, with the energy becoming more accurate as the box size $L$ and the momentum cutoff $k_{c}$ are increased.
The results shown here use a coupling strength of $\alpha=2$. For different $\alpha$ the overall behavior is the same, with the errors caused by finite $L$ and $k_{c}$ being roughly proportional to the ground state energy.
Keeping $L$ constant, the ground state energy decreases with increasing momentum cutoff $k_{c}$ consistent with algebraic convergence, as shown in Fig. 4. The best fit curve is
$E_{0}$ $/(\hbar\omega_{\mathrm{LO}})=-2.3455\pm 0.0039$
$+\left[(0.4110\pm 0.0070)\frac{k_{c}l_{0}}{\pi}\right]^{(-1.0568\pm 0.0451)}.$ (17)
The exponent was found to be close to 1-1 for other values of $\alpha$ as well.
Figure 4: (a) The ground state energy for $\alpha=2$ with $L=6l_{0}$ as a function of the cutoff $k_{c}$. (b) The difference between the ground state energy and the extrapolated energy at infinite $k_{c}$, on a log-log plot. The trend suggests an algebraic convergence of the energy, with an error in the energy proportional to approximately $k_{c}^{-1}$.
Using this, we can estimate that the error incurred in the ground state energy from limiting the maximum phonon wavenumber to $k_{c}=12\pi/l_{0}$ is $(0.0297\pm 0.0064)\hbar\omega_{\mathrm{LO}}$, about $1.3\%$ of the energy. For $k_{c}=4\pi/l_{0}$ the error is $(0.095\pm 0.0082)\hbar\omega_{\mathrm{LO}}$, about $4.1\%$ of the energy.
Figure 5: (a) The ground state energy for $\alpha=2$ with $k_{c}=10\pi/l_{0}$, for various box sizes $L$. (b) The difference between the ground state energy and the extrapolated energy at infinite $L$, on a semi-log plot. In this case, the data imply an exponential convergence of the energy.
Figure 5 suggests an exponential convergence of the energy in the box size $L$ when $k_{c}$ is kept constant. The best fit curve is
$E_{0}$ $/(\hbar\omega_{\mathrm{LO}})=-2.3089\pm 0.0008$
$-(9.5077\pm 0.2381)\exp\left[-(1.5976\pm 0.0132)\frac{L}{l_{0}}\right].$ (18)
Note that the value at infinite $L$ is too high since the momentum cutoff is finite. The error caused by the finite box size at $L=6l_{0}$ can be estimated from the observed exponential convergence to be $(0.00065\pm 0.0011)\hbar\omega_{\mathrm{LO}}$, consistent with no error.
The calculations reported in Fig. 1 were performed at $L=6l_{0}$ and $k_{c}=12\pi/l_{0}$. These calculations with $M=73$ modes used $2\times 10^{5}$ walkers to probe a Hilbert space with dimension $256^{M}\approx 6\times 10^{175}$. The number of non-zero vector coefficients was strongly dependent on $\alpha$. For $\alpha=0.25$ there were about $1.4\times 10^{4}$ non-zero coefficients, whereas for $\alpha=2$ there were about $1.8\times 10^{5}$, almost as many as the number of walkers. Evolving for about $10^{6}$ steps resulted in a stochastic error in the energy estimate of approximately $3\times 10^{-3}\hbar\omega_{\mathrm{LO}}$ for $\alpha=2$. This is an order of magnitude smaller than the estimated discretization error due to the finite value of $k_{c}$.
Appendix B Excited State Energy Convergence
B.1 Bound Excited State
At $\alpha=2$, the first excited state is discrete corresponding to an excitation of the phonon cloud that is bound to the electron (polaron). The energy of this state converges in $k_{c}$ in a similar way to the ground state indicating algebraic convergence. See Fig. 6(a). The best fit curve (at $L=6l_{0}$) is
$E_{1}$ $/(\hbar\omega_{\mathrm{LO}})=-1.4082\pm 0.0046$
$+\left[(0.4429\pm 0.0046)\frac{k_{c}l_{0}}{\pi}\right]^{-1.2002\pm 0.0576}.$ (19)
Figure 6: (a) Energy of the first excited state for $\alpha=2$ and $L=6l_{0}$, with varying momentum cutoff. There is an apparent algebraic convergence, similar to that of the ground state energy. (b) The energy difference between the bound excited state and the ground state for $\alpha=2$ and $L=6l_{0}$, with varying $k_{c}$. The energy difference converges more quickly than the energies of the states separately.
Although the excited state energy has a similarly slow convergence with $k_{c}$ as the ground state, we find that the difference between the two, i.e. the excitation energy, converges much more quickly in $k_{c}$, see Fig. 6(b). Therefore, even though there is a large error in the energies of the ground state and first excited state separately with a momentum cutoff of $k_{c}=4\pi/l_{0}$, this $k_{c}$ is large enough to find the excitation energy. With a momentum cutoff this small, the energies of the ground state and the excited states are too high by about $0.1\hbar\omega_{\mathrm{LO}}$, but the differences between the states are constrained mostly to the low momentum modes resulting in more precise excitation energies. While the numerical data is not accurate enough to distinguish between algebraic and exponential convergence, both extrapolations estimate a systematic error at $k_{c}=4\pi/l_{0}$ that is smaller than the stochastic error bar and consistent with no error.
Figure 7: (a) Energy of the first excited state for $\alpha=2$ and $k_{c}=4\pi/l_{0}$, with varying box size. The rate of convergence is unclear, but even with a slow algebraic convergence the box size error is on the order of $0.01\hbar\omega_{\mathrm{LO}}$. (b) The excitation energy of the bound state with the same parameters. This is not monotonic because the energies of the two states converge differently.
On the other hand, Fig. 7 shows that the energy of the first excited state converges more slowly in $L$ than the ground state, which was already converged at $L=6l_{0}$. With an exponential fit, the exponent is $(-0.54\pm 0.05)L/l_{0}$, indicating a slower convergence than the ground state (A).
Assuming instead an algebraic convergence for the worst case scenario, the error in the excitation energy caused by the finite box size at $L=6l_{0}$ estimated from extrapolation is $(0.0100\pm 0.0018)\hbar\omega_{\mathrm{LO}}$.
B.2 Continuum Excited States
The excited states we find in the continuum have excitation energies larger than $1\hbar\omega_{\mathrm{LO}}$ with a $1/L$ energy dependence only because our computational model is discretized. In the limit $L\to\infty$, all excitation energies of the individual discrete states shown in Fig. 2 would converge to $1\hbar\omega_{\mathrm{LO}}$. Note that this means that the linear interpolation used to find the crossover point for the bound excited state (see Fig. 2 inset) is only valid up to a certain precision. To find the crossing point more precisely with large $L$, one would need to take more data at smaller intervals of $\alpha$.
Appendix C Overcoming the sign problem in excited state FCIQMC
Figure 8: Overcoming the sign problem by walker annihilation. The mean of the shift energy $\langle S\rangle$ is shown as a function of the number of walkers $N_{\mathrm{w}}$ for the ground and first excited states at two values of the interaction strength $\alpha=2.0$ (green) and $\alpha=2.75$ (red), with $L=6l_{0}$ and $k_{c}=4\pi/l_{0}$. The critical number of walkers $N_{c}$ above which the mean shift becomes an estimator for the excited state energy depends on the interaction strength $\alpha$.
The Hamiltonian of the Fröhlich polaron of Eq. (IV) is stoquastic, i.e. all non-zero off-diagonal elements are negative. Thus, it avoids the sign problem for the ground state, as stochastic spawns never change the sign of a walker  [^1] Specifically, all matrix elements of the operator $1-\text{d}\tau(\hat{H}_{\mathrm{LLP}}-S)$ that defines the FCIQMC algorithm are non-negative for small enough $\text{d}\tau$, and a ground state vector with a uniform sign is obtained upon iteration.. This changes for the calculation of excited states where the Gram-Schmidt projector introduces sign-changing contributions to coefficients. In the absence of walker annihilation or at a small number of walkers where annihilation is inefficient, the iterations converge to a ground state of a different Hamiltonian with a lower (incorrect) ground state energy [30]. This leads to an incorrect value of the mean shift energy $\langle S\rangle$. Typically, the presence of a sign problem in Projector Monte Carlo requires a suppression mechanism that introduces uncontrolled approximations (e.g. fixed nodes [74], a constrained path [75], the phaseless [76] or initiator method [77]). Formulated in a discrete Hilbert space, FCIQMC can overcome the sign problem by walker annihilation, avoiding such approximations, if the coefficient vector is sufficiently strongly localized [19, 30, 31].
Figure 8 shows how the mean shift energy depends on the number of walkers for the ground and the first excited state of the Fröhlich polaron problem. For the ground state, the mean shift only carries small stochastic fluctuations and is a valid estimator for the ground state energy regardless of the walker number because the sign problem is absent. The population control bias, which raises $\langle S\rangle$ above the correct value for small walker numbers [24], is smaller than the stochastic errors in this case. For the excited states, the strong dependence of the shift energy on the walker number below a critical value indicates the presence of a sign problem. Above the critical walker number $N_{\mathrm{c}}$, the mean of the shift fluctuates around the correct value of the excited state energy. This critical number is found to be strongly dependent on the value of the coupling strength $\alpha$. We find values of around $N_{\mathrm{c}}\approx 1\times 10^{4}$ for $\alpha=2.0$ and $N_{\mathrm{c}}\approx 1\times 10^{5}$ for $\alpha=2.75$, as seen from Fig. 8. The calculations reported in Fig. 2 and Fig. 3 were obtained with $N_{\mathrm{w}}$ on the order of $10^{6}$ or $10^{7}$, at least an order of magnitude higher than the observed critical values $N_{\mathrm{c}}$ for overcoming the sign problem.
Appendix D Phonon Density
The bound excited state can be studied more closely by looking at the density of virtual phonons in each mode, as shown in Fig. 9. The ground state gains more phonons as the coupling strength increases, with a larger density in the lower momentum modes. The continuum threshold state $|\mathrm{C}\rangle$ has increased phonon density only in the $k=0$ mode, as expected.
Figure 9: The density of virtual phonons in each mode $\langle j|a_{k}^{\dagger}a_{k}|j\rangle$ for the ground state $|j\rangle=|\mathrm{GS}\rangle$ and first two excited states, with $L=6l_{0}$ and $k_{c}=4\pi/l_{0}$, for various coupling strengths. The state $|\mathrm{C}\rangle$ is at the start of the continuum, and always has an excitation energy of $1\hbar\omega_{\mathrm{LO}}$. The state $|\mathrm{B}\rangle$ is a bound excited state for $\alpha\gtrsim 1.73$ and otherwise part of the discretized continuum. For the total phonon numbers see Table 1.
For weak coupling, the second excited state is part of the discretized continuum and mostly a combination of the Fock state with one phonon in each of the $k=\pm 2\pi/L$ modes (one of the $a=2$, $b=0$ states), and the two Fock states with one phonon, in each of those modes respectively (the $a=1$, $b=\pm 1$ states). So for weak coupling the state can be approximated as $\left[c_{1}\left(\hat{a}^{\dagger}_{2\pi/L}+\hat{a}^{\dagger}_{-2\pi/L}\right)+c_{2}\hat{a}^{\dagger}_{2\pi/L}\hat{a}^{\dagger}_{-2\pi/L}\right]|\text{vac}\rangle$ where $|\text{vac}\rangle$ is the phonon vacuum, for some $c_{1}$ and $c_{2}$ depending on $L$. As the coupling strength increases, there is a larger occupation in the other modes. For $\alpha=2$, when this state drops below the continuum to become a bound excited state, this simple description no longer applies. The bound state phonon density then becomes much more weakly affected by the mode discretization and the box size.
The number of phonons in each of the states in Fig. 9 is presented in Table 1.
$\alpha$ $N_{|\mathrm{GS}\rangle}$ $N_{|\mathrm{C}\rangle}$ $N_{|\mathrm{B}\rangle}$
0.5 0.29 1.29 1.95
1.0 0.67 1.67 2.36
1.5 1.19 2.19 2.96
2.0 1.94 2.94 3.68
Table 1: The total number of phonons in the dressing cloud as the integral of the phonon density for the states presented in Fig. 9. The continuum threshold state $|\mathrm{C}\rangle$ contains one more phonon than the ground state. In the strong coupling limit $\alpha\to\infty$, the first bound excited state $|\mathrm{B}\rangle$ has 0.6475 more phonons than the ground state.